### v1

@@ 5.2 GPRFormer Accuracy
Table 2 shows RMSE computed across the cumulative displacement determined from the GPRFormer predictions for each test sequence in the CMU-GPR dataset and the MarsLGPR dataset. It is important to note that the wheel encoder data in the CMU-GPR dataset is very accurate, as demonstrated by the strong quantitative results. This is because the dataset collection occurred on concrete surfaces, where slip was minimal and there were no slopes. The Mars analog terrain in MarsLGPR proves more challenging for wheel encoder odometry estimation. GPRFormer demonstrates improvement over wheel encoder odometry for the more challenging MarsLGPR dataset.
Although [33] does report MSE numbers on the same sequences for the CMU-GPR dataset, we are unable to directly compare with their results as code is not available to replicate their results, and there are not enough details on how their MSE metric was computed to ensure fair comparison. Instead, we include cumulative displacement plots on the CMU-GPR dataset in Figure 8, comparing the ground truth, wheel encoder, and GPRFormer displacement, in order to allow for qualitative comparison with the cumulative displacement plots reported in [33]. Note that during the training process, we train on overlapping windows of GPR data. During test time, we similarly feed in overlapping windows of GPR a-scans as they are provided in real-time from the dataset. In order to account for this potential “double-counting” of displacement from multiple overlapping windows, we take a weighted averaging approach, where the number of predictions from all overlapping windows at each time step is used to produce an average displacement estimate from multiple network predictions.
The RMSE metrics in Table 2 show that for the MarsLGPR dataset, we outperform the wheel encoder baseline by around 30% to 50%. For the CMU-GPR dataset, the wheel encoder baseline has very little noise or drift due to the dataset being collected on flat concrete ground. We do not outperform this non-noisy wheel encoder baseline, but we do show qualitatively good tracking in Figure 8.
The righthand column of Figure 8 shows the individual GPRFormer predictions at each timestep compared with the displacement from the ground truth and wheel encoders. This allows us to inspect whether the network is able to learn local trends and patterns from the GPR data. We see that the network is successfully able to determine the velocity increases and decreases at the start and tail end of each sequence. Some of the small variations in displacement are not always captured, but this may also be due to noise in the ground truth. Since we are measuring displacement within a very small interval, any effects of noise will be amplified.
@@
Figure 9 shows qualitative results for the GPRFormer predictions on MarsLGPR test sequences. Note how the wheel encoder data is much less accurate compared to the RTAB-Map baseline as the wheel encoder data is in the CMU-GPR dataset, which makes it a much less reliable source of odometry. Our GPRFormer network visibly outperforms the wheel encoder displacement as compared to the reference.
@@ 5.3 EKF State Estimation Accuracy
We also demonstrate the resulting full state estimation results when integrating the GPRFormer predictions along with IMU yaw angle into an EKF to estimate $x$ and $y$ position. Table 3 shows a comparison of the EKF with and without the GPRFormer integrated in. We calculate the RMSE Absolute Trajectory Error (ATE) metric on both trajectories in relation to the ground truth after aligning them in the yaw axis.
@@ 6 CONCLUSION & FUTURE WORK
In this work, we proposed a novel framework to leverage GPR sensors aboard planetary rovers to aid in localization. Through our proposed network architecture, we demonstrated the ability for GPRFormer to learn relative displacement from GPR measurements on various types of Martian analog rugged terrain. Additionally, we integrated our prediction network into an EKF sensor fusion framework in order to provide efficient and accurate robot pose estimation. Our GPRFormer EKF outperforms the wheel encoder + IMU only filter, especially in areas of high wheel slippage. We will publicly release our MarsLGPR dataset and code to enable further work in GPR localization for Mars analog environments.
One limitation of the filtering-based framework that we use in this work is the need to carefully tune covariance matrices, which represent uncertainty from GPRFormer. Future work would include an investigation into quantifying the uncertainty from the GPRFormer model in order to provide dynamic and intelligent covariance matrices for uncertainty propagation in the filter.
@@ ACKNOWLEDGMENT
The authors would like to thank the mission support crew at MDRS and the Mars Society, namely Ben Stanley, for their support during field operations. Additionally, the authors would like to thank Madelyn Hoying, Rebecca McCallin, Ben Kazimer, Anna Tretiakova, and Nicole Chan for their assistance in data collection in the field.

### v2

@@ 5.1 GPRFormer Implementation Details
5.2 Baseline Methods
As no code is available for the Wickramanayake [34] baseline, we have re-implemented this method from scratch in order to use it as a comparison. We differentiate our re-implemented version of this baseline as Wick.*. We directly follow the parameters provided in the paper, with a learning rate of 0.0001, an MSE loss function, the Adam optimizer, a batch size of 100, and the same train/test split. We trained the model for 250 epochs on the CMU-GPR dataset and 60 epochs on the MarsLGPR dataset, after which the MSE loss was no longer reduced. The paper reports batch MSE values of 0.266, 0.403, 0.115, and 0.143, our replicated version of the CNN-LSTM had batch MSE values of 0.370, 0.226, 0.123, 0.105 for Sequences 1, 3, 7, and 9, respectively.
@@ 5.3 GPRFormer Accuracy
We evaluate model performance using the Root Mean Squared Error (RMSE) of displacement estimates, reported in millimeters. RMSE is given by:
$\text{RMSE}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}},$ (16)
where $y_{t}$ is the ground truth displacement, $\hat{y}_{t}$ is the predicted displacement, and $T$ is the prediction window. For GPRFormer, $T=k=10$, and for Wick.*, $T=2$. Since the two models generate predictions in overlapping windows of length $T$, we apply an overlap-add accumulation procedure to obtain a consistent per-timestep estimate: each window’s prediction is distributed evenly across its span and averaged with other overlapping contributions, producing a single displacement estimate per timestep. This binning ensures that different prediction granularities are made directly comparable, and also makes the predictions more robust to outlier results from the model.
Table 2 shows RMSE computed across the cumulative displacement for each test sequence in the CMU-GPR dataset and the MarsLGPR dataset. It is important to note that the wheel encoder data in the CMU-GPR dataset is very accurate, as demonstrated by the strong quantitative results. This is because the dataset collection occurred on concrete surfaces, where slip was minimal and there were no slopes. The Mars analog terrain in MarsLGPR proves more challenging for wheel encoder odometry estimation. GPRFormer demonstrates improvement over wheel encoder odometry for the more challenging MarsLGPR dataset. Our method, GPRFormer, outperforms the re-implemented Wick.* method on all eight test sequences across both datasets.
Figure 9 displays qualitative results from the RTAB-Map reference, the wheel encoder baseline, the Wick.* method baseline, and our model GPRFormer on the CMU-GPR dataset. The righthand column shows the individual GPRFormer predictions at each timestep compared with the displacement from the ground truth and wheel encoders. This allows us to inspect whether the network is able to learn local trends and patterns from the GPR data. The Wick.* predictions are very noisy, which is reflected in a higher RMSE metric in Table 2. The lefthand column shows the cumulative displacement over the trajectory, with the direction of the wheel encoder (positive or negative) indicating the direction of movement. Note that in all eight plots, Wick.* has no results on the last section of the trajectory because the method relies on building up 202 consecutive GPR A-scans before making a displacement prediction.
@@
Figure 10 shows qualitative results for GPRFormer and Wick.* on MarsLGPR test sequences. Note how the wheel encoder in the MarsLGPR sequences is much less accurate than the wheel encoder in the CMU-GPR dataset, which makes it a much less reliable source of odometry. Our GPRFormer network visibly outperforms the wheel encoder displacement as compared to the RTAB-Map reference, and this is also reflected quantitatively in Table 2. We note that in Sequence 12-17-01-44, there were not enough GPR traces for the Wick.* method to make any predictions.
5.4 GPRFormer Runtime Performance
In this experiment, we discuss the ability of the GPRFormer model to operate in realtime. We present Table 4, which shows the runtime per sample in seconds for our method and baseline Wick.* [34]. We show runtime for inference on both GPU and CPU, as GPUs are typically inaccessible in space environments. Both methods have very fast runtimes, on the order of hundreds of Hz, although our method outperforms the Wick.* model.
Although the model runtimes are very fast, the most critical bottleneck for performance is the GPR data rate. For the MarsLGPR dataset, GPR data is collected at 1.67 Hz, meaning that our model (which requires 10 consecutive samples) could begin inference after 6 seconds, and then every half second after that. The Wickramanayake method requires two patches of 200 traces, with two traces in between, for a total of 202 traces (over 120 seconds of data collection) in order to begin inference.
Table 4: Performance metrics (memory and runtime) for Wick.* and GPRFormer model inference.
Model Avg. inference runtime (s) Avg. memory (GB)
CPU GPU CPU GPU
Wick.* [34] 0.0706 0.0640 1.23 2.66
Ours 0.0034 0.0032 1.90 1.39
@@ 5.5 EKF State Estimation Accuracy
We also demonstrate the resulting full state estimation results when integrating the GPRFormer predictions along with IMU yaw angle into an EKF to estimate $x$ and $y$ position. As GPRFormer is only able to predict 1D displacement, we increase the covariance of the GPR network predictions when the robot is rotating, as determined by the wheel encoder directions. Table 3 shows a comparison of the EKF with and without the GPRFormer integrated in. We calculate the RMSE Absolute Trajectory Error (ATE) metric on both trajectories in relation to the ground truth after aligning them in the yaw axis.
@@
5.6 Ablation Studies
In this experiment, we investigate how the different components of the network impact the overall model performance. We train five ablated versions of the model on the MarsLGPR dataset, with two additional experiments to further understand the $k$ and $\alpha$ parameters. See Table 5 for a comparison of the performance of each ablation. The metric used in this experiment is RMSE in millimeters across all four MDRS test sequences, where lower is better.
Table 5: Ablation study on GPRFormer network with results in RMSE ($m$) across all four MarsLGPR test sequences. Bold is best.
B-scan filter encoder dropout seq pooling # layers RMSE ($\downarrow$)
15.57
16.58
16.94
16.51
18.25
20.37
5.6.1 B-scan filter pre-processing
This pre-processing step helps make features in the B-scan more distinct and visible. We see in the ablation Table 5 that not including this pre-processing step does make the results of the network worse, but not drastically so. This is likely because the network is still able to learn to find the finer-grained features in the B-scan even without the filtering. However, the inclusion of B-scan filtering likely makes the model learn more efficiently, resulting in an improvement when filtering is used.
5.6.2 Linear encoder
Similarly, removing the linear encoder from the GPRFormer architecture has a moderate effect on the overall network performance, slightly increasing the RMSE. This experiment demonstrates that the B-scan representation of the data is already quite compact and structured, which allows the model to learn well despite having no encoder. However, adding a linear encoder layer does allow for improved results by letting the model learn a more efficient latent space.
5.6.3 Dropout
Dropout is an important technique for improving network generalization and avoiding overfitting. In this variation, we reduce the dropout rate from 0.1 to 0 in order to understand the effects of this component in the architecture, which causes a moderate drop in network performance. The worse GPRFormer performance without the dropout module is likely caused by overfitting on the training set that occurs without this regularization.
5.6.4 Dual sequence pooling
As discussed in the technical approach of the paper, dual sequence pooling is a useful technique to allow the GPRFormer model to benefit from both the pre-transformer encoding and the post-transformer encoding by blending attention weights. In this experiment, we completely remove this from the network, which causes the network to strictly use the transformer encoder latent space. This greatly affects the performance of the model, with the average cumulative RMSE increasing to over 18 mm. This helps confirm that the inclusion of this attention weight pooling step, which balances between raw patterns and transformer-encoded patterns in the data, is a key component of the GPRFormer network.
[…]
Figure 12: A comparison of different $\alpha$ (dual sequence pooling weight) values and their impact on the model accuracy, as measured in RMSE. Note that a $\alpha=0.7$ has the best results.
Additionally, we include a further experiment on the optimal $\alpha$ value, which is the weighting parameter between the pre-transformer and post-transformer features. The $\alpha$ is a learned term, but for this experiment we fixed it to a static value in order to understand the impact of the term on the model performance. We see in Figure 12 that when $\alpha$ is at a value of 0.7–which weights the post-transformer features a little higher than pre-transformer features–it has the best performance in terms of RMSE. When $\alpha$ is unfrozen during normal training, it converges to 0.70102, matching the results of this experiment.
551010151520202525303035354040161617171818kkRMSE ($m$)Impact of $k$ on model accuracy
Figure 13: A comparison of different $k$ (B-scan length) values and their impact on the model accuracy, as measured in RMSE. Note that a $k=10$ has the best results.
5.6.5 Multi-layer transformer encoder
In this experiment, we drop the number of layers and attention heads to just one, and we see this has the most drastic effect on the network performance out of all of the ablations, with an increase in RMSE to over 20 mm. The transformer encoder is one of the core components of the network, and it has enough complexity that it can compensate for some of the other ablation experiments (removing the B-scan filtering step, for example). However, when this block in the network is simplified too far, it loses much of its ability to perform well.
5.6.6 B-scan length k
In this experiment, we experimentally validate our choice to use a $k$ (B-scan length) of 10. In Figure 13, there is a clear minimum RMSE at $k=10$, which is the value we used in our network design. As discussed in the paper sensor model, GPR has a very wide beam width, resulting in spatial overlap with adjacent A-scans. Stacking several traces together provides enough contextual information to the model about subsurface spatial features that the model is able to predict displacement. However, for the GPRFormer architecture the tradeoff between a larger B-scan and better performance starts to degrade after $k=10$. The larger input makes it more difficult for the network to isolate meaningful features.
@@ 6 CONCLUSION & FUTURE WORK
In this work, we proposed a novel framework to leverage GPR sensors aboard planetary rovers to aid in localization. Through our proposed network architecture, we demonstrated the ability for GPRFormer to learn relative displacement from GPR measurements on various types of Martian analog rugged terrain. We compare our method to both a wheel encoder and learned baseline [34], and GPRFormer achieves superior results. Additionally, we integrate our prediction network into an EKF sensor fusion framework in order to provide efficient and accurate robot pose estimation. Our GPRFormer EKF outperforms the wheel encoder + IMU only filter, especially in areas of high wheel slippage. We will publicly release our MarsLGPR dataset and GPRFormer code to enable further work in GPR localization for Mars analog environments.
One limitation of the filtering-based framework that we use in this work is the need to carefully tune covariance matrices, which represent uncertainty from GPRFormer. Future work would include an investigation into quantifying the uncertainty from the GPRFormer model in order to provide dynamic and intelligent covariance matrices for uncertainty propagation in the filter. Additionally, we note that recent advances in 3D neural representations with novel sensors, such as NeRFs with Synthetic Aperture Radar [44] as well as 3D Gaussian Splatting with radar [45] and forward-looking sonar [46], motivate further work in GPR feature maps for localization.
@@ ACKNOWLEDGMENT
The authors would like to thank the mission support crew at MDRS and the Mars Society, namely Ben Stanley, for their support during field operations. Additionally, the authors would like to thank Madelyn MacRobbie, Rebecca McCallin, Ben Kazimer, Anna Tretiakova, and Wing Lam Chan for their assistance in conducting field data collection.